### v1

# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, heterogeneous, homogeneous, and one-to-one, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks. Empirical results demonstrate significant performance improvements in clients’ SLMs with the aid of the LLM. Furthermore, the LLM optimized by FedMKT achieves a performance comparable to that achieved through direct fine-tuning based on clients’ data, highlighting the effectiveness and adaptability of FedMKT.
1 Introduction
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022) and GPT-4 OpenAI (2023), with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. This trend signifies a growing emphasis on developing and utilizing large-scale pre-train models, underscoring their increasing importance in advancing the frontiers of AI. The remarkable success of LLMs in general NLP tasks has also fueled the release of numerous high-performing open-source LLMs, including LLaMA Touvron et al. (2023), ChatGLM Du et al. (2022), OPT Zhang et al. (2022a), igniting a wave of enthusiasm for further exploration and application.
Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023) Fan et al. (2023). The primary challenges include:
  • Domain-Specific Knowledge Privacy. Downstream clients are unable to access the LLM parameters when the LLM is black-box. Hence, they must transmit their domain-specific data to the server that owns the LLM owners for fine-tuning purposes. This process inevitably compromises the privacy of clients’ sensitive, domain-specific data.
  • Constrained Computing Resources. Downstream clients may face limitations in computing resources, which can hinder their ability to fine-tune downloaded LLMs, even if they have access to the server’s LLM. The fine-tuning process requires substantial computational and storage resources, posing a significant barrier for small and medium-sized companies with limited resources to adopt LLMs. Consequently, these companies, which act as clients in FL, are often compelled to fine-tune smaller language models (SLMs) using their domain-specific data, which may result in inferior model performance.
  • Mutual Knowledge Transfer Between the LLM and SLMs. The optimization of both the LLM and SLMs across different parties establishes a positive feedback loop, enabling the continuous evolution of both the LLM and SLMs. Initially, the LLM on the server can disseminate general knowledge and capabilities to the clients’ SLMs. Subsequently, these domain-specific SLMs can return their industry-specific knowledge to the server’s LLM, fostering a collaborative learning environment. However, a significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To address these challenges, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation techniques, FedMKT facilitates effective mutual knowledge transfer between clients’ small language models (SLMs) and a large language model (LLM) located at the server.
As illustrated in Figure 1, FedMKT deploys an LLM on the server and a set of $K$ heterogeneous SLMs across various clients. The cornerstone of FedMKT lies in its selective mutual knowledge transfer process. During each round of federated learning, the clients transmit the output logits of their updated SLMs on the public dataset to the server. The server then selectively aggregates and distills the knowledge encoded within these SLMs output logits into the server-side LLM. This process allows the server LLM to incorporate the domain-specific knowledge learned by the clients, thereby enhancing its comprehensive capabilities. Simultaneously, the server-side LLM also selectively distills its knowledge to the clients’ SLMs, which is similar to the knowledge transfer from clients to the server. By leveraging the knowledge of the server LLM, the clients’ SLMs are able to improve their performance and generalize better to unseen data. To address the model heterogeneity between the LLM and SLMs, FedMKT incorporates a token alignment technique utilizing minimum edit distance (MinED) prior to knowledge transfer. This alignment ensures seamless integration and efficient knowledge transfer between LLM and SLMs.
Extensive experiments conducted on various public LLM and SLMs, including GPT-2 Radford et al. (2019), OPT Zhang et al. (2022a), Bloom Scao et al. (2022), LLaMa Touvron et al. (2023), demonstrate the competitive performance of our FedMKT framework across a range of NLP text generation tasks. In our comprehensive evaluation, we test our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. The results show that the SLMs can achieve significant enhancements with the support of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ domain data directly. Overall, FedMKT represents a practical and effective solution for grounding LLMs in real-world applications while maintaining data privacy protections.
Our contributions are summarized as follows:
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between large language models (LLMs) deployed on the server and small language models (SLMs) residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
  • Selective Knowledge Transfer and Token Alignment. FedMKT implements a selective knowledge transfer mechanism that selectively distills knowledge from the most informative SLMs to the server’s LLM and vice versa. Furthermore, it incorporates a token alignment technique using minimum edit distance (MinED) to address model heterogeneity between LLM and SLMs, ensuring efficient knowledge transfer.
  • Empirical Evaluation and Performance Enhancement. Through extensive experiments on various public LLMs and SLMs across NLP text generation tasks, FedMKT demonstrates significant performance enhancements in clients’ SLMs with the aid of the server’s LLM. Additionally, the LLM optimized by FedMKT achieves comparable performance to direct fine-tuning on clients’ data, highlighting the effectiveness and adaptability of the proposed framework for grounding LLMs in real-world applications while maintaining data privacy protections.
Figure 1: FedMKT (Federated Mutual Knowledge Transfer for Large and Small Language Models Framework.)
2 Related Work
2.1 Model Heterogeneous Federated Learning
Federated learning has emerged as a promising approach for collaborative machine learning across multiple devices while preserving data privacy. However, traditional federated learning frameworks often assume homogeneity among client models, which may not hold in real-world scenarios due to varying computational capabilities, network conditions, and data distributions. This limitation has led to the development of model heterogeneous federated learning (MHFL), which aims to address the challenges associated with heterogeneity. Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models.
Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Each client then calculates the distance between the global logits and the local logits of their local heterogeneous model on the same public data sample, using this as a distillation loss for training the local model.
Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. These methods integrate mutual learning into personalized federated learning, aiming to enhance the performance of both the global and personalized heterogeneous models.
Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
While previous work has mainly focused on computer vision (CV) scenarios, there has been limited exploration of MHFL in large language model (LLM). This gap in the literature motivates our work, which aims to explore MHFL in the context of LLMs.
2.2 Federated Learning for LLMs
Adapting large language models (LLMs) Yang et al. (2023); Zhou et al. (2023); OpenAI (2022, 2023); Touvron et al. (2023) to downstream tasks typically involve the comprehensive fine-tuning of all model parameters, a process known as full fine-tuning. However, this approach can be prohibitively expensive, particularly for domain-specific tasks. To address this challenge, Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) have been proposed. These methods efficiently adapt LLMs to specific domains or tasks by fine-tuning only a subset of model parameters while keeping the rest frozen.
PEFT methods offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for large language models (LLMs). A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM Fan et al. (2023). Specifically, the FedPETuning methods proposed by Zhang et al. (2022b) have demonstrated a significant reduction in communication overhead, achieving a reduction of 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting. Additionally, they found that PETuning methods can effectively reduce local model adaptation costs for clients in FL systems. These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods. These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
3 The Proposed Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer for large and small language models framework. This framework encompasses fine-tuning scenarios in an FL setting where an LLM is run on the server while SLMs are distributed among the clients. A notable aspect of this architecture is the heterogeneity in designs among the client SLMs and their contrast with the server LLM. We begin by defining the specific problem addressed in this study, followed by a detailed introduction to our approach. Finally, we delve into the computational and communication complexities and the privacy-preserving aspects of our FedMKT framework.
3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. We assume that the $K$ clients execute the same text generation task, but they hold heterogeneous SLM $g_{\phi_{k}}$.
The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without ever disclosing any private data. The threat model we consider is a standard semi-honest security model, wherein a semi-honest server(S) attempts to deduce the private data belonging to the client(c). Specifically,
  • Each client possesses its own local private dataset $\mathcal{D}^{k}$ and owns a common public dataset $D^{p}$. Each client $k$ aims to collectively train its local private SLM $g_{\phi_{k}}$ without divulging their private data. Subsequently, each client computes the output logits on the public dataset and securely uploads them to the server. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (1)
  • The server also owns the common public dataset $D^{p}$, and receives the output logits of different clients’ SLMs on the public dataset. Subsequently, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on the public dataset. The objective can be formulated as follows:
    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}^{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
    The server aims to transfer knowledge from the clients’ SLMs $g_{\phi_{k}}$ to its owned LLM $f_{\psi}$.
  • Once the knowledge transfer from SLMs to LLM is completed on the server, the server dispatches the LLM’s output logits on the public dataset $D^{p}$ to all the clients. Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on the public dataset $D^{p}$. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}^{p},{\psi})$ (3)
    The clients aim to transfer knowledge from LLM $f_{\psi}$ to SLMs $g_{\phi_{k}}$ to enhance SLMs in the clients.
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We elaborate on FedMKT in Section 3.2.
3.2 FedMKT
FedMKT represents an innovative and parameter-efficient framework for federated mutual knowledge transfer between large and small language models. Its primary objective is to foster the collaborative advancement of both server-side LLM and client-side SLMs. Additionally, in our work, we specifically focus on text-generation tasks. We illustrate the FedMKT in Figure 1 and describe the associated training algorithm in Algorithm 1.
Algorithm 1 FedMKT
1:
2:$K$: number of clients;
3:$T$: total number of rounds;
4:$\theta^{0}_{k}$: the initialized global adapter of SLM $g_{\phi_{k}+\theta_{k}}$;
5:$\omega^{0}$: the initialized adapter of LLM $f_{\psi+\omega}$;
6:$\eta_{\omega}$: the learning rate for optimizing LLM $f_{\psi+\omega}$;
7:$\eta_{\theta}$: the learning rate for optimizing SLM $g_{\phi_{k}+\theta_{k}}$.
8:$f_{\psi+\omega}$, $g_{\phi_{1}+\theta_{1}}$,$g_{\phi_{2}+\theta_{2}}$,…,$g_{\phi_{K}+\theta_{K}}$.
9:for $t$ in communication round $T$ do
[…]
11:     $p_{k}^{t},l_{k}^{t}\leftarrow\textbf{ClientUpdate1}(t)$
12:     $\triangleright$ selective knowledge transfer from SLMs to LLM based on $\mathcal{D}^{p}$
13:     Token Alignment from SLMs to LLM on $\mathcal{D}^{p}$.
14:     for each epoch $r\in[R]$ do
15:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$
16:     end for
17:     $\omega^{t+1}=\omega^{t,R}$
18:     $p^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
19:     $l^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
[…]
21:
[…]
23:     for each client $k$ (in parallel) do
24:         $\triangleright$ local fine-tuning based on $\mathcal{D}^{k}$
25:         for each local epoch $e\in[E]$ do
26:              $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}^{k}$
27:         end for
28:         $p_{k}^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
29:         $l_{k}^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
30:     end for
31:
32:     //ClientUpdate2($p^{t},l^{t}$):
33:     for each client $k$ (in parallel) do
34:         Token Alignment from LLM to SLM $g_{\phi_{k}+\theta_{k}}$ on $\mathcal{D}^{p}$.
35:         $\triangleright$ selective knowledge transfer from LLM to SLMs based on $\mathcal{D}^{p}$
36:         for each local epoch $e\in[E,2E]$ do
37:              $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$
38:         end for
39:         $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$
40:     end for
41:end for
Bidirectional Token Alignment with MinED. A notable challenge that arises during the alignment of output logits distributions is the mismatch between the tokenizers of different LLMs and SLMs, such as the Bloom tokenizer and the LLaMa tokenizer. Consider the sentence, ”we utilize the dynamic programming approach to align tokens” as an example. Utilizing the Bloom tokenizer would segment it into the following tokens: [’we’, ’utilize’, ’the,’ ’dynamic,’ ’programming,’ ’approach,’ ’to,’ ’align,’ ’tokens’]. However, if the LLaMa tokenizer were used, the segmentation would be: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. This misalignment, coupled with model heterogeneity, poses a significant hurdle. Therefore, achieving token alignment across multiple SLMs and LLM is paramount for effective knowledge transfer in a federated learning setting. Proper alignment guarantees the accurate mapping of output logits and probabilistic distribution matrices.
To tackle this issue, we utilize dynamic programming techniques that foster resilient alignment and minimize the impact of tokenizer mismatches, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Given that tokens produced by distinct tokenizers for the same sequence often exhibit minimal disparities, we employ a minimum edit distance (MinED) strategy. This approach maps tokens from various tokenizers based on their MinED, allowing for flexible token alignment. By relaxing the alignment criteria, we retain significant information within the distribution matrices while introducing negligible errors. This method ensures effective knowledge transfer in the federated learning setting, overcoming the challenges posed by tokenizer mismatches and model heterogeneity.
Knowledge transfer between LLMs and SLMs involves a bidirectional token alignment process, which is divided into two stages.
  • In the first stage, clients transfer knowledge of their SLMs to the LLM on the server, aiming to enhance the LLM’s capacity. During this phase, the SLMs generate output logits that are uploaded to the server. Here, the tokens produced by the SLMs serve as the source, while the tokens from the LLM serve as the target. It is crucial to accurately align the SLM tokens with the corresponding LLM tokens.
  • In the second stage, the LLM on the server transfers its knowledge back to the clients’ SLMs, aiming to improve their performance. As the LLM’s output logits are downloaded to the clients, the roles reverse: the LLM’s tokens become the source, and the SLM tokens become the target.
This bidirectional alignment ensures seamless integration and effective knowledge transfer between the models. We can address the challenges posed by tokenizer mismatches and model heterogeneity by leveraging this approach.
Selective Mutual Knowledge Transfer Between LLM and SLMs. Each client $k$ inserts a local small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote client $k$ local SLM with the added $\theta_{k}$ as $g_{\phi_{k}+\theta_{k}}$. Each client $k$ then trains its local LoRA adapter module $\theta_{k}$ using its private dataset. Consequently, Eq.(1) can be reformulated as follows:
$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (4)
$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}^{k}}$ $\ell_{\text{TA}}^{k}(g_{\phi_{k}+\theta_{k}}(x),y)$
where $\ell_{\text{TA}}$ is the task loss for training the local LoRA module $\theta_{k}$ in the client $k$. The original model parameter $\phi_{k}$ of client $k$’s local SLM is frozen during training.
The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. The server conducts the knowledge transfer from each local SLM $g_{\phi_{k}+\theta_{k}}$ to the LLM $f_{\psi+\omega}$ and through supervised fine-tuning and selective knowledge distillation based on the public dataset $\mathcal{D}^{p}$.
We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
where $\ell_{\text{CE}}$ is the cross-entropy loss; the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
The selective mutual knowledge distillation losses for $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ models (denoted as $\mathcal{L}_{\text{KD}}^{f}$ and $\mathcal{L}_{\text{KD}}^{g}$) are formulated as follows:
$\mathcal{L}^{f}_{\text{KD}}(\omega;\mathcal{D}^{p},\theta_{1},\theta_{2},...,\theta_{K})$ (7)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(f_{\psi+\omega}(x),g_{\phi_{k}+\theta_{k}}(x))$
$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\mathcal{D}^{p},\psi+\omega)$ (8)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),f_{\psi+\omega}(x))$
Since clients’ knowledge may adversely affect LLM on the server side in heterogeneous environments, and the reverse is also true, where the server-side LLM could potentially have a negative influence on the clients’ SLMs, it is essential to evaluate the prediction quality of both LLM and SLMs for each public data sample. To facilitate positive knowledge transfer, we propose selective strategies on both the server and client sides, termed DualMinCE. DualMinCE work as follows: For each public sample, we compute the cross-entropy loss across all SLMs and LLM, comparing the prediction distribution with the gold labels. Subsequently, when knowledge needs to be transferred from SLMs to the LLM, we select the SLM exhibiting the minimum cross-entropy loss to effectively distill its knowledge into the LLM residing on the server. Conversely, when transferring knowledge from the LLM to SLMs, we only proceed with the distillation if the LLM’s cross-entropy loss is lower than that of the SLMs. This ensures that the most accurate and informative knowledge is distilled between the LLM and SLMs.
Combining Eq.(5) and Eq.(7), we reformulate the knowledge transfer from SLMs to LLM conducted on the server to enhance LLM as follows:
$\mathcal{L}_{2}$ $=\lambda\mathcal{L}_{\text{FT}}^{f}+(1-\lambda)\mathcal{L}_{\text{KD}}^{f}$ (9)
Combining Eq.(6) and Eq.(8), we reformulate the knowledge transfer from LLM to SLMs conducted on the clients to enhance SLMs as follows:
$\mathcal{L}_{3}$ $=\frac{1}{K}\sum_{k=1}^{K}(\lambda\mathcal{L}_{\text{FT}}^{g}+(1-\lambda)\mathcal{L}_{\text{KD}}^{g})$ (10)
where $\lambda$ is the hyperparameter that control the weight of mutual knowledge transfer.
FedMKT Workflow. FedMKT cultivates a mutually beneficial relationship between the server and clients, fostering a symbiosis where both parties reap the rewards of collective knowledge and expertise embedded in their respective language models. By harnessing the distinct strengths of server-side LLM and client-side SLMs, FedMKT pioneers a new frontier in efficient and effective federated learning within the realm of LLM. This innovative approach enables the collaborative evolution of LLM and SLMs. The workflow of FedMKT is outlined as follows:
  1. 1.
    In the $t$-th communication round, the $K$ clients train their respective LoRA adapters using their private data. This step allows the clients to adapt their models to their specific data distributions.
  2. 2.
    After local training, the $K$ clients upload their SLMs’ output logits and the corresponding cross-entropy losses computed on the public dataset to the server.
  3. 3.
    On the server side, token alignment is performed from the SLMs to the LLM, guaranteeing compatibility between the SLMs and the LLM.
  4. 4.
    On the server side, knowledge is seamlessly selectively transferred from the SLMs to the LLM, where the server utilizes the SLMs’ output logits and cross-entropy losses to enrich the LLM’s understanding.
  5. 5.
    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes the LLM’s output logits and the corresponding cross-entropy losses of the public dataset.
  6. 6.
    The server disseminates the LLM’s output logits and the corresponding cross-entropy losses of the public dataset to all the clients.
  7. 7.
    On the client side, the token alignment flow reverses, and token alignment is performed from the LLM to SLMs.
  8. 8.
    On the client side, the knowledge transfer flow reverses, with the LLM’s insights being selectively transferred to the SLMs, where the clients leverage the LLM’s output logits and cross-entropy losses to further enhance their SLMs.
3.3 Computation and Communication Complexity
One of the key advantages of FedMKT is its computational efficiency. By leveraging PEFT, the framework significantly reduces the number of parameters that need to be updated during fine-tuning. For instance, it consumes just 0.12% of the computational cost associated with fine-tuning all parameters in OPT-1.3B when using FedMKT. This leads to faster training times and reduced computational requirements, making it more feasible to fine-tune LLM and SLMs in a federated learning setting.
In terms of communication complexity, FedMKT minimizes the amount of data exchanged between clients and the server. Instead of transmitting entire models(For example, OPT-1.3B is about 1.3B floating-point numbers), clients only share the output logits and corresponding cross-entropy losses of the public dataset with the server. Suppose there are $N=1000$ public text samples with a text sequence length of $S=512$ and a top token size of $K=16$. The communication cost, denoted as $Cost_{com}$, would be calculated as follows: $Cost_{com}=N*S*K=1000*512*16=8M$ floating-point numbers. This approach reduces communication overhead, allowing for more efficient data transmission and enhancing scalability in federated learning scenarios.
Overall, the computational efficiency and reduced communication complexity of FedMKT make it a promising framework for federated mutual knowledge transfer for large and small Language models, particularly in scenarios where computational resources and communication bandwidth are limited.
Setting Server Client-1 Client-2 Client-3 Client-4
Heterogeneous LLaMa2- 7B GPT-2-xlarge(1.5B) OPT-1.3B Bloom-1.1B LLaMa2-1.3B
Homogeneous LLaMa2- 7B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B
Homogeneous LLaMa2- 7B OPT-1.3B OPT-1.3B OPT-1.3B OPT-1.3B
One-to-One LLaMa2- 7B - - - LLaMa2-1.3B
One-to-One LLaMa2- 7B - OPT-1.3B - -
Table 1: FedMKT: LLM and SLMs Setting
Task Method GPT-2-xlarge OPT-1.3B Bloom-1.3B LLaMa2-1.3B
RTE SLM-ZS 52.4 52.7 52.7 49.8
Standalone 65.7 62.5 58.1 55.6
FedMKT-SLM 70.4 65.7 61.7 58.8
WIC SLM-ZS 49.8 50.8 50 50
Standalone 59.3 52.1 59.1 50.7
FedMKT-SLM 63.2 62.2 61.1 51.9
BoolQ SLM-ZS 61.3 58.4 59.0 61.0
Standalone 71.1 74.1 69.9 69.7
FedMKT-SLM 75.1 76.8 71.4 75.1
CQA SLM-ZS 36.7 41.9 33.8 30.1
Standalone 56.0 58.6 44.7 56.7
FedMKT-SLM 58.3 60.5 50.8 57.0
OpenBookQA SLM-ZS 22.4 23.4 19.6 23.2
Standalone 27.4 26.6 21.0 25.8
FedMKT-SLM 27.8 28.0 23.2 27.2
ARC-E SLM-ZS 58.3 57.0 51.5 53.1
Standalone 59.3 57.9 56.9 60.3
FedMKT-SLM 59.8 59.6 57.5 60.8
ARC-C SLM-ZS 25.0 23.4 23.6 26.7
Standalone 28.2 28.4 24.9 28.5
FedMKT-SLM 30.2 29.4 26.6 30.0
S-NI SLM-ZS 5.0 5.2 5.1 5.8
Standalone 27.9 26.1 10.6 33.4
FedMKT-SLM 34.2 36.0 15.1 37.3
DialogSum SLM-ZS 5.4 6.4 4.9 5.7
Standalone 22.3 19.8 13.1 21.4
FedMKT-SLM 23.2 20.9 14.9 21.6
Table 2: Heterogeneous Setting SLMs Performance Comparison: FedMKT versus Baseline.
4 Experiments
4.1 Setup
We set up a scenario involving three clients and one server to evaluate the FedMKT framework using various LLMs and SLMs.
Models. We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al. (2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al. (2019), OPT-1.3B Zhang et al. (2022a), Bloom-1.1B Scao et al. (2022) and LLaMa2-1.3B Xia et al. (2023). In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. Table 1 details the model configurations for both the LLM and SLMs in the distinct scenarios.
Datasets. We comprehensively evaluate FedMKT on 7 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CommonsenseQA(CQA) Talmor et al. (2018), OpenBookQA Mihaylov et al. (2018), ARC-E Clark et al. (2018), ARC-CClark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
Baselines. We conduct a comparative analysis of our FedMKT framework against several baselines to evaluate its performance. These baselines included:
  • LLM-ZS, which represents the zero-shot (ZS) capabilities of the large language model (LLM) on the server;
  • SLM-ZS, which represents the zero-shot performance of the small language model (SLM) on the clients;
  • Standalone, where each client independently fine-tunes its local model using its own private dataset;
  • FedAvg, which represents the standard federated averaging algorithm and is only used in the SLM comparisons of homogeneous settings.
  • LLM2SLM, where the LLM is not updated in the FedMKT and is LLM only used to transfer knowledge to SLM in the FedMKT. It is only used in the SLM comparisons of the One-to-One setting.
  • Centralized-LLM, where the server’s LLM is fine-tuned locally using the entirety of the private datasets combined with a public distillation dataset.
Evaluation Metrics. We evaluate the performance of fine-tuned LLM and SLMs on both QA datasets and instruction-following datasets. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L. This approach allows us to comprehensively assess the models’ capabilities across different tasks, providing a clear comparison of their performance.
4.2 Heterogeneous Setting: Performance Evaluations
Tables 2 and 3 demonstrate the performance comparisons of our approach against other baselines in the heterogeneous setting. Specifically, Table 2 focuses on the performance of SLMs, while Table 3 highlights the performance of LLM.
As illustrated in Table 2, FedMKT-SLM demonstrates notably superior performance in comparison to both SLM-ZS and Standalone models. For instance, in the RTE QA dataset, FedMKT-SLM demonstrates significant performance gains compared to both SLM-ZS and Standalone models across various SLMs. Specifically, when using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 34% and Standalone by 7%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 25% and Standalone by 5%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 17% improvement over SLM-ZS and a 6% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 18% and Standalone by 6%. In the DialogSum instruction-following dataset, FedMKT-SLM again demonstrates impressive results. When using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 330% and Standalone by 4%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 227% and Standalone by 6%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 204% improvement over SLM-ZS and a 14% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 279% and Standalone by 0.9%. This highlights that within our framework, SLMs are able to effectively leverage the knowledge transferred from LLM, leading to enhanced model capabilities.
As illustrated in Table 3, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM. For instance, in the RTE QA dataset, FedMKT-LLM outperforms LLM-ZS by 30% and achieves a performance level that is nearly on par with Centralized-LLM, reaching approximately 96% of its fine-tuning performance. Similarly, in the DialogSum instruction-following dataset, FedMKT-LLM outperforms LLM-ZS by a significant 185% and achieves approximately 90% of the fine-tuning performance of Centralized-LLM. This significant achievement signifies that FedMKT-LLM effectively facilitates the acquisition of knowledge from all clients by the server. This process enables the LLM to be fine-tuned utilizing client data without necessitating access to the entire model. Moreover, the clients’ raw private data is not required to be uploaded to the LLM server, thus preserving privacy while enabling the LLM to leverage the collective knowledge of all clients.
Task Method LLaMa2-7B
RTE LLM-ZS 63.1
Centralized-LLM 85.9
FedMKT-LLM 82.3
WIC LLM-ZS 50.3
Centralized-LLM 70.4
FedMKT-LLM 61.3
BoolQ LLM-ZS 70.1
Centralized-LLM 87.6
FedMKT-LLM 84.6
CQA LLM-ZS 39.5
Centralized-LLM 69.5
FedMKT-LLM 71.8
OpenBookQA LLM-FS 31.8
Centralized-LLM 39.2
FedMKT-LLM 37.8
ARC-E LLM-ZS 69.3
Centralized-LLM 76.9
FedMKT-LLM 76.1
ARC-C LLM-ZS 40.0
Centralized-LLM 48.9
FedMKT-LLM 44.7
S-NI LLM-ZS 12.0
Centralized-LLM 49.3
FedMKT-LLM 41.4
DialogSum LLM-ZS 8.5
Centralized-LLM 27.7
FedMKT-LLM 24.2
Table 3: Heterogeneous Setting LLM Performance Comparison: FedMKT versus Baseline.
4.3 Homogeneous Setting: Performance Evaluations
In the homogeneous setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. It’s important to note that the number of clients is four, and within each client, all the SLMs is homogeneous. In this setting, to facilitate a comparison with FedAvg, we incorporated a FedAvg process involving all four clients within the FedMTK-SLM framework subsequent to mutual knowledge transfer. Furthermore, it’s worth noting that the standalone score here represents the average score obtained from the four clients working independently.
Tables 4 demonstrate the performance comparisons of our approach against other baselines in the homogeneous setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance than the Standalone and FedAvg approaches due to the assistance provided by LLM. Meanwhile, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM.
Task Method LLaMa2 OPT
CQA SLM-ZS 30.1 41.9
Standalone 56.4 58.1
FedAvg 56.4 58.6
FedMKT-SLM-FedAvg 58.6 61.5
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM-FedAvg 68.8 71.3
ARC-C SLM-ZS 26.7 23.4
Standalone 30.4 28.5
FedAvg 29.7 28.6
FedMKT-SLM-FedAvg 31.7 29.9
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM-FedAvg 46.2 46.2
ARC-E SLM-ZS 53.1 57.0
Standalone 60.3 58.8
FedAvg 60.6 59.0
FedMKT-SLM-FedAvg 61.7 60.0
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.9
FedMKT-LLM-FedAvg 74.9 75.8
Table 4: Homogeneous Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
4.4 One-to-One Setting: Performance Evaluations
In the One-to-One setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. Noth that both configurations involve only one client.
Tables 5 demonstrate the performance comparisons of our approach against other baselines in the one-to-one setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance compared to both the Standalone and LLM2SLM approaches. Compared to Standalone, FedMTK-SLM achieves superior results due to the assistance provided by LLM. When compared to LLM2SLM, FedMTK-SLM further improves performance by updating the LLM, which incorporates knowledge from client-side SLM. These findings underscore the benefits of federated mutual knowledge transfer between LLMs and SLMs. Meanwhile, compared to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses LLMs’ zero-shot performance and rivals Centralized-LLM’s fine-tuning performance.
Task Method LLaMa2 OPT
RTE SLM-ZS 49.8 52.7
Standalone 55.6 62.5
LLM2SLM 57.0 64.3
FedMKT-SLM 61.0 63.2
LLM-ZS 63.1 63.1
Centralized-LLM 85.9 85.9
FedMKT-LLM 82.3 79.1
CQA SLM-ZS 30.1 41.9
Standalone 56.7 58.6
LLM2SLM 56.76 59.1
FedMKT-SLM 56.84 60.7
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM 69.0 71.0
ARC-C SLM-ZS 26.7 23.4
Standalone 30.3 28.8
LLM2SLM 30.1 29.6
FedMKT-SLM 30.8 30.4
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM 45.9 45.8
ARC-E SLM-ZS 53.1 57.0
Standalone 57.0 57.9
LLM2SLM 60.7 59.1
FedMKT-SLM 60.8 59.7
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.1
FedMKT-LLM 74.8 75.8
Table 5: One-to-One Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
5 Conclusions and Future Work
In this work, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models. FedMKT bridges the gap between the server-side large language model (LLM) and clients’ small language models (SLMs), enabling bidirectional knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios — Heterogeneous, Homogeneous, and One-to-One — we have demonstrated the effectiveness of FedMKT By utilizing a range of publicly available LLMs and SLMs on various NLP text generation tasks, we’ve shown that FedMKT simultaneously boosts the performance of both LLMs and SLMs. The results highlight the framework’s adaptability and practicality, paving the way for integrating LLMs into real-world applications while protecting data privacy. As large language models continue to transform the landscape of artificial intelligence, FedMKT offers a promising approach for their responsible and effective utilization.
Limitations and Future works. Our current study is constrained by computational and storage resources, preventing us from exploring larger model sizes. We plan to address this in future research.

### v2

@@
 We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance.
-Through extensive experiments across three distinct scenarios, heterogeneous, homogeneous, and one-to-one, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks.
-Empirical results demonstrate significant performance improvements in clients’ SLMs with the aid of the LLM.
-Furthermore, the LLM optimized by FedMKT achieves a performance comparable to that achieved through direct fine-tuning based on clients’ data, highlighting the effectiveness and adaptability of FedMKT.
+Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks.
+Empirical results demonstrate that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
+FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models
+Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
 1 Introduction
 Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities.
-The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022) and GPT-4 OpenAI (2023), with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks.
-This trend signifies a growing emphasis on developing and utilizing large-scale pre-train models, underscoring their increasing importance in advancing the frontiers of AI.
-The remarkable success of LLMs in general NLP tasks has also fueled the release of numerous high-performing open-source LLMs, including LLaMA Touvron et al.
-(2023), ChatGLM Du et al.
-(2022), OPT Zhang et al.
-(2022a), igniting a wave of enthusiasm for further exploration and application.
+The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), and LLaMA Touvron et al.
+(2023) with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks.
 Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al.
@@
 (2023).
-The primary challenges include:
-  • Domain-Specific Knowledge Privacy.
-Downstream clients are unable to access the LLM parameters when the LLM is black-box.
-Hence, they must transmit their domain-specific data to the server that owns the LLM owners for fine-tuning purposes.
-This process inevitably compromises the privacy of clients’ sensitive, domain-specific data.
-  • Constrained Computing Resources.
-Downstream clients may face limitations in computing resources, which can hinder their ability to fine-tune downloaded LLMs, even if they have access to the server’s LLM.
-The fine-tuning process requires substantial computational and storage resources, posing a significant barrier for small and medium-sized companies with limited resources to adopt LLMs.
-Consequently, these companies, which act as clients in FL, are often compelled to fine-tune smaller language models (SLMs) using their domain-specific data, which may result in inferior model performance.
-  • Mutual Knowledge Transfer Between the LLM and SLMs.
-The optimization of both the LLM and SLMs across different parties establishes a positive feedback loop, enabling the continuous evolution of both the LLM and SLMs.
-Initially, the LLM on the server can disseminate general knowledge and capabilities to the clients’ SLMs.
-Subsequently, these domain-specific SLMs can return their industry-specific knowledge to the server’s LLM, fostering a collaborative learning environment.
-However, a significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits.
+The primary challenges include domain-specific knowledge Privacy, constrained computing resources, and mutual knowledge transfer between the LLM and SLMs.
+A significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits.
 The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle.
 Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
-To address these challenges, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models.
-By leveraging the complementary strengths of federated learning and knowledge distillation techniques, FedMKT facilitates effective mutual knowledge transfer between clients’ small language models (SLMs) and a large language model (LLM) located at the server.
+To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models.
+By leveraging the complementary strengths of federated learning and knowledge distillation, FedMKT facilitates effective mutual knowledge transfer between clients’ SLMs and the LLM owned by the server.
 As illustrated in Figure 1, FedMKT deploys an LLM on the server and a set of $K$ heterogeneous SLMs across various clients.
@@
 This alignment ensures seamless integration and efficient knowledge transfer between LLM and SLMs.
-Extensive experiments conducted on various public LLM and SLMs, including GPT-2 Radford et al.
-(2019), OPT Zhang et al.
-(2022a), Bloom Scao et al.
-(2022), LLaMa Touvron et al.
-(2023), demonstrate the competitive performance of our FedMKT framework across a range of NLP text generation tasks.
-In our comprehensive evaluation, we test our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One.
-The results show that the SLMs can achieve significant enhancements with the support of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ domain data directly.
-Overall, FedMKT represents a practical and effective solution for grounding LLMs in real-world applications while maintaining data privacy protections.
 Our contributions are summarized as follows:
   • Federated Mutual Knowledge Transfer Framework.
-FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between large language models (LLMs) deployed on the server and small language models (SLMs) residing on clients.
+FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between an LLM deployed on the server and SLMs residing on clients.
 This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
@@
   • Empirical Evaluation and Performance Enhancement.
-Through extensive experiments on various public LLMs and SLMs across NLP text generation tasks, FedMKT demonstrates significant performance enhancements in clients’ SLMs with the aid of the server’s LLM.
-Additionally, the LLM optimized by FedMKT achieves comparable performance to direct fine-tuning on clients’ data, highlighting the effectiveness and adaptability of the proposed framework for grounding LLMs in real-world applications while maintaining data privacy protections.
-Figure 1: FedMKT (Federated Mutual Knowledge Transfer for Large and Small Language Models Framework.)
+Extensive experiments conducted based on various publicly available LLMs and SLMs demonstrate the competitive performance of FedMKT across a wide range of NLP text-generation tasks.
+We evaluate FedMKT with heterogeneous, Homogeneous, and One-to-One settings.
+The results show that the performance of SLMs can be significantly enhanced with the help of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ data centralized.
+Figure 1: Overview of the proposed FedMKT workflow. Each communication round of FedMKT involves 11 steps to fine-tune the server’ LLM and clients’ SLMs.
 2 Related Work
 2.1 Model Heterogeneous Federated Learning
-Federated learning has emerged as a promising approach for collaborative machine learning across multiple devices while preserving data privacy.
-However, traditional federated learning frameworks often assume homogeneity among client models, which may not hold in real-world scenarios due to varying computational capabilities, network conditions, and data distributions.
-This limitation has led to the development of model heterogeneous federated learning (MHFL), which aims to address the challenges associated with heterogeneity.
+Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning.
 Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures.
@@
 (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits.
-Each client then calculates the distance between the global logits and the local logits of their local heterogeneous model on the same public data sample, using this as a distillation loss for training the local model.
 Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al.
@@
 (2023), design a small homogeneous model and a large heterogeneous model in each client.
-These methods integrate mutual learning into personalized federated learning, aiming to enhance the performance of both the global and personalized heterogeneous models.
 Split learning-based MHFL approaches, such as FedClassAvg Jang et al.
@@
 (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
-While previous work has mainly focused on computer vision (CV) scenarios, there has been limited exploration of MHFL in large language model (LLM).
-This gap in the literature motivates our work, which aims to explore MHFL in the context of LLMs.
+While previous works have mainly focused on computer vision scenarios, the literature has limitedly explored MHFL in LLMs.
+This gap motivates this study, which aims to explore MHFL in the context of LLMs.
 2.2 Federated Learning for LLMs
-Adapting large language models (LLMs) Yang et al.
-(2023); Zhou et al.
-(2023); OpenAI (2022, 2023); Touvron et al.
-(2023) to downstream tasks typically involve the comprehensive fine-tuning of all model parameters, a process known as full fine-tuning.
-However, this approach can be prohibitively expensive, particularly for domain-specific tasks.
-To address this challenge, Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al.
+Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al.
 (2019); He et al.
@@
 (2021); Li and Liang (2021); Hu et al.
-(2021) have been proposed.
-These methods efficiently adapt LLMs to specific domains or tasks by fine-tuning only a subset of model parameters while keeping the rest frozen.
-PEFT methods offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for large language models (LLMs).
+(2021) offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for LLMs.
 A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al.
@@
 (2023).
-Specifically, the FedPETuning methods proposed by Zhang et al.
-(2022b) have demonstrated a significant reduction in communication overhead, achieving a reduction of 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting.
-Additionally, they found that PETuning methods can effectively reduce local model adaptation costs for clients in FL systems.
+For example, the FedPETuning Zhang et al.
+(2022b) has demonstrated a significant reduction in communication overhead, reducing 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting.
 These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods.
@@
 By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
-3 The Proposed Method
-In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer for large and small language models framework.
-This framework encompasses fine-tuning scenarios in an FL setting where an LLM is run on the server while SLMs are distributed among the clients.
-A notable aspect of this architecture is the heterogeneity in designs among the client SLMs and their contrast with the server LLM.
-We begin by defining the specific problem addressed in this study, followed by a detailed introduction to our approach.
-Finally, we delve into the computational and communication complexities and the privacy-preserving aspects of our FedMKT framework.
+3 The Proposed FedMKT Method
+In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer approach for large and small language models.
+The FedMKT primarily comprises two key modules: Bidirectional Token Alignment and Selective Mutual Knowledge Transfer.
+We will elaborate on these two modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
 3.1 Problem Definition
 We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$.
-We assume that the $K$ clients execute the same text generation task, but they hold heterogeneous SLM $g_{\phi_{k}}$.
-The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without ever disclosing any private data.
-The threat model we consider is a standard semi-honest security model, wherein a semi-honest server(S) attempts to deduce the private data belonging to the client(c).
-Specifically,
-  • Each client possesses its own local private dataset $\mathcal{D}^{k}$ and owns a common public dataset $D^{p}$.
-Each client $k$ aims to collectively train its local private SLM $g_{\phi_{k}}$ without divulging their private data.
-Subsequently, each client computes the output logits on the public dataset and securely uploads them to the server.
+Each client owns a local private dataset $\mathcal{D}_{k}$ with $N$ training samples, and all clients and the server share a public dataset $\mathcal{D}_{p}$.
+The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without disclosing any private data.
+We assume that the $K$ clients execute the same text generation task, but they may hold heterogeneous or homogeneous SLM models.
+The collaboration between clients and the server involves the following sub-procedures:
+  • Each client $k$ trains its SLM $g_{\phi_{k}}$ using its private data $\mathcal{D}_{k}$. The objective is formulated as follows:
+    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}_{k}\}_{k=1}^{K})$ (1)
+  • Each client computes the output logits on $\mathcal{D}_{p}$ and securely uploads them to the server.
+Upon receiving output logits of all clients, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on $\mathcal{D}_{p}$.
 The objective can be formulated as follows:
-    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (1)
-  • The server also owns the common public dataset $D^{p}$, and receives the output logits of different clients’ SLMs on the public dataset.
-Subsequently, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on the public dataset.
+    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}_{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
+    The server aims to transfer knowledge from the clients’ SLMs $g_{\phi_{k}}$ to its owned LLM $f_{\psi}$.
+  • The server dispatches the LLM’s output logits on $\mathcal{D}_{p}$ to all the clients.
+Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on $\mathcal{D}_{p}$.
 The objective can be formulated as follows:
-    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}^{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
-    The server aims to transfer knowledge from the clients’ SLMs $g_{\phi_{k}}$ to its owned LLM $f_{\psi}$.
-  • Once the knowledge transfer from SLMs to LLM is completed on the server, the server dispatches the LLM’s output logits on the public dataset $D^{p}$ to all the clients.
-Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on the public dataset $D^{p}$.
-The objective can be formulated as follows:
-    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}^{p},{\psi})$ (3)
-    The clients aim to transfer knowledge from LLM $f_{\psi}$ to SLMs $g_{\phi_{k}}$ to enhance SLMs in the clients.
-FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We elaborate on FedMKT in Section 3.2.
-3.2 FedMKT
-FedMKT represents an innovative and parameter-efficient framework for federated mutual knowledge transfer between large and small language models.
-Its primary objective is to foster the collaborative advancement of both server-side LLM and client-side SLMs.
-Additionally, in our work, we specifically focus on text-generation tasks.
-We illustrate the FedMKT in Figure 1 and describe the associated training algorithm in Algorithm 1.
+    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}_{p},{\psi})$ (3)
+    The clients aim to transfer knowledge from LLM $f_{\psi}$ to enhance their SLMs.
+We consider the server semi-honest, meaning that the server may try to recover the private data of clients from the information it observes.
+FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner.
+We illustrate the workflow of FedMKT in Figure 1 and elaborate on the associated training algorithm in Algorithm 1.
 Algorithm 1 FedMKT
@@
 2:$K$: number of clients;
-3:$T$: total number of rounds;
-4:$\theta^{0}_{k}$: the initialized global adapter of SLM $g_{\phi_{k}+\theta_{k}}$;
-5:$\omega^{0}$: the initialized adapter of LLM $f_{\psi+\omega}$;
-6:$\eta_{\omega}$: the learning rate for optimizing LLM $f_{\psi+\omega}$;
-7:$\eta_{\theta}$: the learning rate for optimizing SLM $g_{\phi_{k}+\theta_{k}}$.
+3:$T$: total number of communication rounds;
+4:$R$: local number of rounds in the server;
+5:$E$: local number of rounds in the client;
+6:$\eta_{\omega}$: the learning rate of LLM $f_{\psi+\omega}$;
+7:$\eta_{\theta}$: the learning rate of SLM $g_{\phi_{k}+\theta_{k}}$.
 8:$f_{\psi+\omega}$, $g_{\phi_{1}+\theta_{1}}$,$g_{\phi_{2}+\theta_{2}}$,…,$g_{\phi_{K}+\theta_{K}}$.
-9:for $t$ in communication round $T$ do
+9:// Server side:
+10:for $t$ in communication round $T$ do
 […]
-11:     $p_{k}^{t},l_{k}^{t}\leftarrow\textbf{ClientUpdate1}(t)$
-12:     $\triangleright$ selective knowledge transfer from SLMs to LLM based on $\mathcal{D}^{p}$
-13:     Token Alignment from SLMs to LLM on $\mathcal{D}^{p}$.
-14:     for each epoch $r\in[R]$ do
-15:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$
-16:     end for
-17:     $\omega^{t+1}=\omega^{t,R}$
-18:     $p^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
-19:     $l^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
+12:     Token Alignment from SLMs to LLM.
+13:     $\tilde{\mathcal{S}}_{0}\leftarrow\textbf{DualMinCE}(\mathcal{D}_{p},f_{\psi+\omega},\{\mathcal{S}_{k}\}_{k=1}^{K})$.
+14:     // knowledge transfer based on $\mathcal{D}_{p}$ and $\tilde{\mathcal{S}}_{0}$.
+15:     for each epoch $r\in[R]$ do
+16:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$.
+17:     end for
+18:     $\omega^{t+1}=\omega^{t,R}$.
+19:     Compute $\mathcal{S}_{0}=\{l^{i}_{0},p^{i}_{0}\}_{i=1}^{N}$ based on $\mathcal{D}_{p}$.
 […]
-21:
+21:end for
+22:
+23:ClientUpdate1($t$):
+24:for each client $k$ (in parallel) do
+25:     // local fine-tuning based on $\mathcal{D}_{k}$.
+26:     for each local epoch $e\in[E]$ do
+27:         $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}$.
+28:     end for
+29:     Compute $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ based on $\mathcal{D}_{p}$.
+30:end for
+31:Upload $\{\mathcal{S}_{k}\}_{k=1}^{K}$ to the server
+32:
+33:ClientUpdate2($t,\mathcal{S}_{0}$):
+34:for each client $k$ (in parallel) do
+35:     Token Alignment from LLM to SLMs.
+36:     $\tilde{\mathcal{S}}_{k}\leftarrow$ $\textbf{DualMinCE}(\mathcal{D}_{p}$, $g_{\phi_{k}+\theta_{k}}$, $\mathcal{S}_{0}$).
+37:     // knowledge transfer based on $\mathcal{D}_{p}$ and $\tilde{\mathcal{S}}_{k}$.
+38:     for each local epoch $e\in[E,2E]$ do
+39:         $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$.
+40:     end for
+41:     $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$.
+42:end for
+Algorithm 2 DualMinCE
+1:
+2:$\mathcal{D}_{p}$: the public dataset;
 […]
-23:     for each client $k$ (in parallel) do
-24:         $\triangleright$ local fine-tuning based on $\mathcal{D}^{k}$
-25:         for each local epoch $e\in[E]$ do
-26:              $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}^{k}$
-27:         end for
-28:         $p_{k}^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
-29:         $l_{k}^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
-30:     end for
-31:
-32:     //ClientUpdate2($p^{t},l^{t}$):
-33:     for each client $k$ (in parallel) do
-34:         Token Alignment from LLM to SLM $g_{\phi_{k}+\theta_{k}}$ on $\mathcal{D}^{p}$.
-35:         $\triangleright$ selective knowledge transfer from LLM to SLMs based on $\mathcal{D}^{p}$
-36:         for each local epoch $e\in[E,2E]$ do
-37:              $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$
-38:         end for
-39:         $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$
-40:     end for
-41:end for
-Bidirectional Token Alignment with MinED.
-A notable challenge that arises during the alignment of output logits distributions is the mismatch between the tokenizers of different LLMs and SLMs, such as the Bloom tokenizer and the LLaMa tokenizer.
-Consider the sentence, ”we utilize the dynamic programming approach to align tokens” as an example.
+4:$\mathcal{S}_{k}=\{(l_{k}^{i},p_{k}^{i})\}_{i=1}^{N},k=0\text{ or }\lceil K\rceil$: loss-logit pairs passed from either the server or clients.
+5:$\mathcal{S}$.
+6:$\tilde{\mathcal{S}}\leftarrow\{\}$triangleright initialize an empty set of selective knowledge.
+7:for each $x^{i}$ in $\mathcal{D}_{p}$ do
+8:     $l^{i}_{\text{local}}\leftarrow h(x^{i})$
+9:     $k^{*}=\begin{cases}\arg\min\limits_{k}{(l^{i}_{k})},&\text{if }k=\lceil K\rceil\\ 0,&\text{if }k=0\end{cases}$
+10:     $\tilde{\mathcal{S}}\leftarrow\tilde{\mathcal{S}}+(x^{i},p^{i}_{k^{*}})$ if $l^{i}_{k^{*}}<l^{i}_{\text{local}}$
+11:end for
+3.2 Bidirectional Token Alignment
+A significant challenge in aligning output logits distributions lies in the mismatch between tokenizers of different LLM and SLMs, exemplified by Bloom and LLaMa.
+Consider the sentence, "we utilize the dynamic programming approach to align tokens" as an example.
 Utilizing the Bloom tokenizer would segment it into the following tokens: [’we’, ’utilize’, ’the,’ ’dynamic,’ ’programming,’ ’approach,’ ’to,’ ’align,’ ’tokens’].
 However, if the LLaMa tokenizer were used, the segmentation would be: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’].
-This misalignment, coupled with model heterogeneity, poses a significant hurdle.
-Therefore, achieving token alignment across multiple SLMs and LLM is paramount for effective knowledge transfer in a federated learning setting.
-Proper alignment guarantees the accurate mapping of output logits and probabilistic distribution matrices.
-To tackle this issue, we utilize dynamic programming techniques that foster resilient alignment and minimize the impact of tokenizer mismatches, as evidenced in studies Wan et al.
+To tackle this issue, we adopt dynamic programming techniques to promote robust alignment, as evidenced in studies Wan et al.
 (2024); Fu et al.
 (2023).
-Given that tokens produced by distinct tokenizers for the same sequence often exhibit minimal disparities, we employ a minimum edit distance (MinED) strategy.
-This approach maps tokens from various tokenizers based on their MinED, allowing for flexible token alignment.
-By relaxing the alignment criteria, we retain significant information within the distribution matrices while introducing negligible errors.
-This method ensures effective knowledge transfer in the federated learning setting, overcoming the challenges posed by tokenizer mismatches and model heterogeneity.
-Knowledge transfer between LLMs and SLMs involves a bidirectional token alignment process, which is divided into two stages.
-  • In the first stage, clients transfer knowledge of their SLMs to the LLM on the server, aiming to enhance the LLM’s capacity.
-During this phase, the SLMs generate output logits that are uploaded to the server.
-Here, the tokens produced by the SLMs serve as the source, while the tokens from the LLM serve as the target.
-It is crucial to accurately align the SLM tokens with the corresponding LLM tokens.
-  • In the second stage, the LLM on the server transfers its knowledge back to the clients’ SLMs, aiming to improve their performance.
-As the LLM’s output logits are downloaded to the clients, the roles reverse: the LLM’s tokens become the source, and the SLM tokens become the target.
-This bidirectional alignment ensures seamless integration and effective knowledge transfer between the models.
-We can address the challenges posed by tokenizer mismatches and model heterogeneity by leveraging this approach.
-Selective Mutual Knowledge Transfer Between LLM and SLMs.
-Each client $k$ inserts a local small low-rank adapter parameterized by $\theta_{k}$ into its local SLM.
+Utilizing LLaMa2 and Bloom as illustrative examples, we establish an optimized vocabulary mapping table based on minimum edit distance (MinED).
+This mapping table identifies the closest Bloom token for each LLaMa2 token (e.g., ’utilize’ for ’util’).
+We then tokenize a sentence using both tokenizers and apply a dynamic programming algorithm to determine the optimal matching path.
+When multiple LLaMa2 tokens align to a single Bloom token (e.g., ’util’ and ’ize’ aligning to ’utilize’), we handle them according to the mapping table.
+Please refer to Appendix B for more details.
+In FedMKT, a bidirectional token alignment process occurs before knowledge transfer between LLMs and SLMs.
+One the one hand, when clients transfer knowledge from their SLMs to the server’s LLM, the server aligns SLM tokens to LLM tokens.
+On the other hand, when the server transfers knowledge from its LLM back to clients’ SLMs, each client aligns LLM tokens to its SLM tokens.
+3.3 Selective Mutual Knowledge Transfer Between LLM and SLMs
+To transfer knowledge between the server and clients efficiently, we leverage LoRA to fine-tune the server’s LLM and clients’ SLMs.
+Specifically, each client $k$ inserts a small low-rank adapter parameterized by $\theta_{k}$ into its local SLM.
 We denote client $k$ local SLM with the added $\theta_{k}$ as $g_{\phi_{k}+\theta_{k}}$.
-Each client $k$ then trains its local LoRA adapter module $\theta_{k}$ using its private dataset.
-Consequently, Eq.(1) can be reformulated as follows:
-$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (4)
-$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}^{k}}$ $\ell_{\text{TA}}^{k}(g_{\phi_{k}+\theta_{k}}(x),y)$
-where $\ell_{\text{TA}}$ is the task loss for training the local LoRA module $\theta_{k}$ in the client $k$.
-The original model parameter $\phi_{k}$ of client $k$’s local SLM is frozen during training.
-The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$.
+Likewise, the server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$.
 We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$.
-The server conducts the knowledge transfer from each local SLM $g_{\phi_{k}+\theta_{k}}$ to the LLM $f_{\psi+\omega}$ and through supervised fine-tuning and selective knowledge distillation based on the public dataset $\mathcal{D}^{p}$.
-We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
-$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
-$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
+During the whole federated learning training process, $\theta_{k},k=1,...,K$ and $\omega$ are trained, while $\phi_{k},k=1,...,K$ and $\psi$ are frozen.
+Before transferring knowledge to the server, each client $k$ trains its LoRA adapter $\theta_{k}$ using its private dataset $\mathcal{D}_{k}$. Consequently, Eq.(1) can be reformulated as follows:
+$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}_{k}\}_{k=1}^{K})$ (4)
+$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\ell_{\text{TA}}($ $g_{\phi_{k}+\theta_{k}}(x),y)$
+where $\ell_{\text{TA}}$ is the task loss for training $\theta_{k}$ of each client $k$. The original model parameter $\phi_{k}$ of client $k$’s SLM is frozen during training.
+Then, both the server and clients fine-tune their LoRA adapters based on a shared public dataset $\mathcal{D}_{p}$.
+We formulate the losses of fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
+$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}_{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
+$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}_{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
 where $\ell_{\text{CE}}$ is the cross-entropy loss; the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
-The selective mutual knowledge distillation losses for $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ models (denoted as $\mathcal{L}_{\text{KD}}^{f}$ and $\mathcal{L}_{\text{KD}}^{g}$) are formulated as follows:
-$\mathcal{L}^{f}_{\text{KD}}(\omega;\mathcal{D}^{p},\theta_{1},\theta_{2},...,\theta_{K})$ (7)
-$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(f_{\psi+\omega}(x),g_{\phi_{k}+\theta_{k}}(x))$
-$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\mathcal{D}^{p},\psi+\omega)$ (8)
-$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),f_{\psi+\omega}(x))$
-Since clients’ knowledge may adversely affect LLM on the server side in heterogeneous environments, and the reverse is also true, where the server-side LLM could potentially have a negative influence on the clients’ SLMs, it is essential to evaluate the prediction quality of both LLM and SLMs for each public data sample.
-To facilitate positive knowledge transfer, we propose selective strategies on both the server and client sides, termed DualMinCE.
-DualMinCE work as follows: For each public sample, we compute the cross-entropy loss across all SLMs and LLM, comparing the prediction distribution with the gold labels.
-Subsequently, when knowledge needs to be transferred from SLMs to the LLM, we select the SLM exhibiting the minimum cross-entropy loss to effectively distill its knowledge into the LLM residing on the server.
-Conversely, when transferring knowledge from the LLM to SLMs, we only proceed with the distillation if the LLM’s cross-entropy loss is lower than that of the SLMs.
-This ensures that the most accurate and informative knowledge is distilled between the LLM and SLMs.
+Next, the server and clients conduct selective knowledge transfer to each other.
+The motivation for applying selective knowledge transfer is that some clients’ knowledge may adversely affect the performance of LLM on the server and vice versa in a heterogeneous environment.
+Therefore, it is critical to guarantee that the knowledge transferred between the server and clients is positive to the performance of LLM and SLMs.
+To this end, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
+DualMinCE aims to select knowledge that is positive to the performance of the server’s LLM from clients and vice versa.
+Specifically, when knowledge needs to be transferred from SLMs to the LLM, each client $k$ computes a knowledge set $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ consisting of loss-logit pairs through its local model based on the public dataset $\mathcal{D}_{p}$.
+Then, all $K$ clients send their $\{\mathcal{S}_{k}\}_{k=1}^{K}$ to the server.
+By leveraging DualMinCE (see Algorithm 2 for detail), the server picks a logit $p_{k^{*}}^{i}$ with the smallest loss from $\{l^{i}_{k},p^{i}_{k}\}_{k=1}^{K}$ and adds $p_{k^{*}}^{i}$ to a selective knowledge set $\tilde{\mathcal{S}}_{0}$ if the loss $l_{k^{*}}^{i}$ of $p_{k^{*}}^{i}$ is smaller than the loss $l_{\text{local}}^{i}$ computed through the server’s local LLM based on $x^{i}$ for each $x^{i}$ in $\mathcal{D}_{p}$.
+Next, the server leverages the knowledge distillation loss, denoted as $\mathcal{L}_{\text{KD}}^{f}$, to fine-tune $f_{\psi+\omega}$:
+$\mathcal{L}^{f}_{\text{KD}}(\omega;\tilde{\mathcal{S}}_{0})=$ $\mathbb{E}_{(x,p)\sim\tilde{\mathcal{S}}_{0}}\ell_{\text{CE}}(f_{\psi+\omega}(x),p)$ (7)
+Likewise, each client $k$ leverages DualMinCE to form its selective knowledge set $\tilde{\mathcal{S}}_{k}$ from the knowledge $\mathcal{S}_{0}$ sent from the server.
+Each client $k$ leverages the following knowledge distillation loss to fine-tune its local model $g_{\phi_{k}+\theta_{k}}$:
+$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\tilde{\mathcal{S}}_{k})=$ $\mathbb{E}_{(x,p)\sim\tilde{\mathcal{S}}_{k}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),p)$ (8)
 Combining Eq.(5) and Eq.(7), we reformulate the knowledge transfer from SLMs to LLM conducted on the server to enhance LLM as follows:
@@
 $\mathcal{L}_{3}$ $=\frac{1}{K}\sum_{k=1}^{K}(\lambda\mathcal{L}_{\text{FT}}^{g}+(1-\lambda)\mathcal{L}_{\text{KD}}^{g})$ (10)
-where $\lambda$ is the hyperparameter that control the weight of mutual knowledge transfer.
-FedMKT Workflow.
-FedMKT cultivates a mutually beneficial relationship between the server and clients, fostering a symbiosis where both parties reap the rewards of collective knowledge and expertise embedded in their respective language models.
-By harnessing the distinct strengths of server-side LLM and client-side SLMs, FedMKT pioneers a new frontier in efficient and effective federated learning within the realm of LLM.
-This innovative approach enables the collaborative evolution of LLM and SLMs.
+where $\lambda$ is the hyperparameter that controls the weight of mutual knowledge transfer.
+Setting Server Client-1 Client-2 Client-3 Client-4
+Heterogeneous LLaMa2- 7B GPT-2-xlarge(1.5B) OPT-1.3B Bloom-1.1B LLaMa2-1.3B
+Homogeneous LLaMa2- 7B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B
+Homogeneous LLaMa2- 7B OPT-1.3B OPT-1.3B OPT-1.3B OPT-1.3B
+One-to-One LLaMa2- 7B - - - LLaMa2-1.3B
+One-to-One LLaMa2- 7B - OPT-1.3B - -
+Table 1: The five different settings we utilize to evaluate FedMKT.
+Task Method GPT-2-xlarge OPT-1.3B Bloom-1.1B LLaMa2-1.3B LLaMa2-7B
+RTE Centralized - - - - 85.9
+Zero-Shot 52.4 52.7 52.7 49.8 63.2
+Standalone 65.7 62.5 58.1 55.6 -
+FedMKT 70.4 65.7 61.7 58.8 82.3
+WIC Centralized - - - - 70.4
+Zero-Shot 49.8 50.8 50 50 50.3
+Standalone 59.3 52.2 59.1 50.6 -
+FedMKT 63.2 62.2 61.1 51.9 61.3
+BoolQ Centralized - - - - 87.6
+Zero-Shot 61.3 58.4 59.0 61.0 70.1
+Standalone 71.1 74.1 69.7 69.9 -
+FedMKT 75.1 76.8 71.4 75.1 85.0
+CQA Centralized - - - - 69.5
+Zero-Shot 36.7 41.9 33.8 30.1 39.5
+Standalone 56.0 58.6 44.7 56.7 -
+FedMKT 58.3 60.5 50.8 57.0 71.8
+ARC-E Centralized - - - - 76.9
+Zero-Shot 58.3 57.0 51.5 53.1 69.3
+Standalone 59.3 57.9 56.9 60.4 -
+FedMKT 59.8 59.6 57.5 60.8 76.1
+ARC-C Centralized - - - - 48.9
+Zero-Shot 25.0 23.4 23.6 26.7 40.0
+Standalone 28.2 28.4 24.9 28.5 -
+FedMKT 30.2 29.4 26.6 30.0 44.7
+S-NI Centralized - - - - 49.3
+Zero-Shot 5.0 5.2 5.1 5.8 12.0
+Standalone 27.9 26.1 10.6 33.4 -
+FedMKT 34.2 36.0 15.1 37.3 41.4
+DialogSum Centralized - - - - 27.7
+Zero-Shot 5.4 6.4 4.9 5.7 8.5
+Standalone 22.3 19.8 13.2 21.4 -
+FedMKT 23.2 20.9 14.9 21.6 24.2
+Table 2: Method Performance Comparison in the Heterogeneous setting.
+We evaluate FedMKT with 8 different tasks.
+In all the 8 tasks, the server is deployed with a LLaMa2-7B model, and the 4 clients are deployed with a GPT-2-xlarge, a OPT-1.3B, a Bloom-1.1B, and a LLaMa2-1.3B, respectively.
+The ’-’ indicates a method does not apply to the corresponding participant (either the server or the client).
+4 Experiments
+4.1 Setup
+We set up a federated learning scenario involving four clients and one server to evaluate the FedMKT using various publicly available LLMs and SLMs.
+Models.
+We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al.
+(2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al.
+(2019), OPT-1.3B Zhang et al.
+(2022a), Bloom-1.1B Scao et al.
+(2022) and LLaMa2-1.3B Xia et al.
+(2023).
+In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One.
+Table 1 details the setup for the LLM and SLMs in different settings.
+Datasets.
+We evaluate FedMKT comprehensively on 6 QA datasets and 2 instruction-following datasets.
+Specifically, for QA tasks, we use RTE Wang et al.
+(2019), WTC Wang et al.
+(2019), BoolQ Clark et al.
+(2019), CommonsenseQA(CQA) Talmor et al.
+(2018), ARC-E Clark et al.
+(2018), ARC-C Clark et al.
+(2018) to evaluate FedMKT.
+As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al.
+(2022), DialogSum Chen et al.
+(2021).
+Baselines. We conduct a comparative analysis of FedMKT against the following baselines:
+  • Centralized, in which the server’s LLM is fine-tuned locally using the datasets combining private datasets of involved clients and the public dataset.
+In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are used to fine-tune the LLM;
+  • Zero-Shot, representing the zero-shot capabilities of LLM or SLMs (without fine-tuning);
+  • Standalone, in which each client independently fine-tunes its local SLM using its private dataset;
+  • FedAvg, representing the standard federated averaging algorithm. FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
+  • LLM2SLM, representing FedMKT involving one server with an LLM and one client with an SLM.
+The LLM is not updated and is used to transfer knowledge to SLM.
+LLM2SLM is only used in the One-to-One setting.
+Evaluation Metrics. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
+4.2 Evaluation on Heterogeneous Setting
+In the Heterogeneous setting, the server is deployed with a LLaMa2-7B model, and the 4 clients are deployed with a GPT-2-xlarge, a OPT-1.3B, a Bloom-1.1B, and a LLaMa2-1.3B, respectively.
+Table 2 reports the performance comparisons of FedMKT against baselines on 8 tasks.
+Tables 2 show that FedMKT performs superior over Zero-Shot and Standalone on all clients’ SLMs.
+Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM.
+These empirical results demonstrate that, by leveraging FedMKT, SLMs are able to effectively leverage the knowledge transferred from the LLM, leading to enhanced model capabilities.
+Table 2 also shows that FedMKT outperforms Zero-Shot and Centralized on the LLaMa2-7B of the server.
+For instance, on the RTE QA dataset, FedMKT outperforms Zero-Shot by 30% and achieves a performance level that is nearly on par with Centralized, reaching approximately 96% of its fine-tuning performance.
+This significant achievement signifies that FedMKT effectively facilitates the acquisition of knowledge from all clients by the server.
+4.3 Evaluation on Homogeneous Setting
+We conduct experiments with two Homogeneous settings, as shown in Table 1.
+The first setting (denoted as S1) involves one server-side LLaMa2-7B and four client-side LLaMa2-1.3B.
+The second setting (denoted as S2) involves one server-side LLaMa2-7B and four client-side OPT-1.3B.
+Table 3 reports the performance comparisons of FedMKT against baselines in the two Homogeneous settings.
+The top sub-table and the bottom sub-table compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively.
+The top sub-table of Table 3 shows that FedMKT significantly outperforms Zero-Shot on the server’s LLM (i.e., LLaMa2-7B) in the two Homogeneous settings.
+It also shows that FedMKT achieves comparable performance of the Centralized scenario, in which the server’ LLM is fine-tuned using all clients’ data and the public data combined.
+The bottom sub-table of Table 3 shows that FedMTK performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM.
+For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
+Task Method S1: Server LLaMa2-7B S2: Server LLaMa2-7B
+CQA Zero-Shot 39.5 39.5
+Centralized 69.5 69.5
+FedMKT 68.8 71.3
+ARC-C Zero-Shot 40.0 40.0
+Centralized 49.4 49.4
+FedMKT 46.2 46.2
+ARC-E Zero-Shot 69.3 69.3
+Centralized 75.5 75.5
+FedMKT 74.9 74.8
+Task Method S1: Clients LLaMa2-1.3B S2: Clients OPT-1.3B
+CQA Zero-Shot 30.1 41.9
+Standalone 56.4 58.1
+FedAvg 56.4 58.6
+FedMKT 58.6 61.5
+ARC-C Zero-Shot 26.7 23.4
+Standalone 30.4 28.5
+FedAvg 29.7 28.6
+FedMKT 31.7 29.9
+ARC-E Zero-Shot 53.1 57.0
+Standalone 60.3 57.9
+FedAvg 60.6 58.8
+FedMKT 61.7 60.1
+Table 3: Method Performance Comparison in Homogeneous settings.
+We evaluate FedMKT using two homogeneous settings.
+The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and four client-side LLaMa2-1.3B SLMs, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and four client-side OPT-1.3B SLMs.
+The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively.
+The results reported in the bottom sub-table are the average of all clients.
+4.4 Evaluation on One-to-One Setting
+We evaluate FedMKT using two One-to-One settings.
+The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and one client-side LLaMa2-1.3B SLM, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and one client-side OPT-1.3B SLM.
+Table 4 reports the performance comparisons of FedMKT against baselines in the two One-to-One settings.
+The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively.
+The top sub-table of Table 4 shows that FedMKT notably surpasses Zero-Shot and rivals Centralized on the performance of the server’s LLM.
+The bottom sub-table of Table 4 shows that FedMTK achieves superior SLM performance over Zero-Shot, Standalone, and LLM2SLM due to the assistance of LLM.
+These empirical results demonstrate the effectiveness of FedMKT in transferring knowledge between the LLM and SLMs.
+Task Method S1: Server LLaMa2-7B S2: Server LLaMa2-7B
+CQA Zero-Shot 39.5 39.5
+Centralized 69.0 68.3
+FedMKT 69.0 71.0
+ARC-C Zero-Shot 40.0 40.0
+Centralized 45.9 48.6
+FedMKT 45.9 45.8
+ARC-E Zero-Shot 69.3 69.3
+Centralized 74.4 73.6
+FedMKT 74.8 74.8
+Task Method S1: Clients LLaMa2-1.3B S2: Clients OPT-1.3B
+CQA Zero-Shot 30.1 41.9
+Standalone 56.7 58.6
+LLM2SLM 56.76 59.1
+FedMKT 56.84 60.7
+ARC-C Zero-Shot 26.7 23.4
+Standalone 30.3 28.8
+LLM2SLM 30.1 29.6
+FedMKT 30.8 30.4
+ARC-E Zero-Shot 53.1 57.0
+Standalone 57.0 57.9
+LLM2SLM 60.7 58.4
+FedMKT 60.8 58.5
+Table 4: Method Performance Comparison in One-to-One settings.
+We evaluate FedMKT using two one-to-one settings.
+The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and one client-side LLaMa2-1.3B SLM, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and one client-side OPT-1.3B SLM.
+The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and a client’s SLM, respectively.
+5 Conclusions
+In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models.
+FedMKT bridges the gap between the server-side LLM and clients’ SLM, enabling selective mutual knowledge transfer while preserving data privacy.
+Through extensive experiments across three distinct scenarios, we have demonstrated that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
+Limitations
+In this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations.
+Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters Li and Wang (2019); Cho et al.
+(2022), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data.
+This issue warrants further investigation.
+Furthermore, our study is limited by computational and storage constraints, which preclude the exploration of larger language models.
+This highlights the inherent trade-off between utility and efficiency.
+Our future research aims to investigate and optimize this trade-off.
+Appendix A FedMKT Workflow
 The workflow of FedMKT is outlined as follows:
@@
   2. 2.
-    After local training, the $K$ clients upload their SLMs’ output logits and the corresponding cross-entropy losses computed on the public dataset to the server.
+    After local training, each client $k$ computes a knowledge set $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ consisting of loss-logit pairs through its local model based on the public dataset.
   3. 3.
+    Each client $k$ upload $\mathcal{S}_{k}$ to the server.
+  4. 4.
     On the server side, token alignment is performed from the SLMs to the LLM, guaranteeing compatibility between the SLMs and the LLM.
+  5. 5.
+    On the server side, knowledge is selected from the SLMs to the LLM according to Algorithm 2.
+  6. 6.
+    On the server side, knowledge is transferred from the SLMs to the LLM based on the selected knowledge.
+  7. 7.
+    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes a knowledge set $\mathcal{S}_{0}=\{l^{i}_{0},p^{i}_{0}\}_{i=1}^{N}$ consisting of loss-logit pairs through LLM based on the public dataset.
+  8. 8.
+    The server disseminates $\mathcal{S}_{0}$ to all the clients.
+  9. 9.
+    On the client side, the token alignment flow reverses, and token alignment is performed from the LLM to SLMs.
+  10. 10.
+    On the client side, knowledge is selected from the LLM to each client SLM according to Algorithm 2.
+  11. 11.
+    On the client side, knowledge is transferred from the LLM to each client SLM based on the selected knowledge.
+Appendix B Implementation Details of Token Alignment
+In our work, we engage in a bidirectional token alignment procedure, encompassing the alignment of SLM tokens with their corresponding LLM tokens, and vice versa.
+Both alignments adhere to a similar methodology.
+Presently, we shall elaborate on the process of aligning LLM tokens with their matching SLM tokens.
+To map the predicted token logits from the LLaMa2-7B (LLM) model to the Bloom-1.1B (SLM) model, several steps must be undertaken.
+The detailed process is as follows:
+  1. 1.
+    Building an Optimal Vocabulary Mapping Table:
+    1. (a)
+      For each token in the LLaMa2 vocabulary, iterate through the Bloom vocabulary.
+    2. (b)
+      Use edit distance as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary.
+    3. (c)
+      If there are multiple token with the same minimum edit distance, choose the one with the lexicographically smallest order.
+    4. (d)
+      Save this mapping relationship in the optimal vocabulary mapping table.
+  2. 2.
+    Tokenization and Alignment:
+    1. (a)
+      Tokenize the sentence "we utilize the dynamic programming approach to align tokens" using both the LLaMa2 and Bloom tokenizers.
+    2. (b)
+      To align the two tokenization results and determine the optimal matching path, we utilize a dynamic programming algorithm.
+As an illustration, consider the tokenization outputs from LLaMa2 and Bloom.
+LLaMa2’s tokenization yields: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’].
+In contrast, Bloom’s tokenization produces: [’we’, ’utilize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’].
+In this instance, seven terms from LLaMa2 align perfectly with those from Bloom, such as "we" and "dynamic".
+Notably, the LLaMa2 tokens ’util’ and ’ize’ collectively map to the single Bloom token ’utilize’.
+In scenarios where multiple tokens align to one, like the 2-to-1 case of ’util’ and ’ize’ mapping to ’utilize’, we consider ’utilize’ as a match for ’util’ based on an optimal vocabulary mapping.
+  3. 3.
+    Logits Mapping:
+    1. (a)
+      Iterate through each token $t_{t}$ in the Bloom tokenization result.
+    2. (b)
+      For each $t_{t}$, check if it uniquely matches a token $t_{s}$ in the LLaMa2 tokenization result.
+    3. (c)
+      If $t_{t}$ uniquely matches $t_{s}$, then for each token $t_{p}$ in the Top-$K$ predicted token of $t_{s}$ from LLaMa2 and its corresponding logit $logit_{p}$: Find the position $pos$ in the Bloom vocabulary that corresponds to $t_{p}$ using the optimal vocabulary mapping table.
+If $pos$ has not been assigned a value before, copy $logit_{p}$ to the corresponding position in the Bloom logits distribution matrix $logit_{t}$.
+    4. (d)
+      If $t_{t}$ does not have a unique match, generate one-hot logits for $t_{t}$.
   4. 4.
-    On the server side, knowledge is seamlessly selectively transferred from the SLMs to the LLM, where the server utilizes the SLMs’ output logits and cross-entropy losses to enrich the LLM’s understanding.
-  5. 5.
-    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes the LLM’s output logits and the corresponding cross-entropy losses of the public dataset.
-  6. 6.
-    The server disseminates the LLM’s output logits and the corresponding cross-entropy losses of the public dataset to all the clients.
-  7. 7.
-    On the client side, the token alignment flow reverses, and token alignment is performed from the LLM to SLMs.
-  8. 8.
-    On the client side, the knowledge transfer flow reverses, with the LLM’s insights being selectively transferred to the SLMs, where the clients leverage the LLM’s output logits and cross-entropy losses to further enhance their SLMs.
-3.3 Computation and Communication Complexity
+    Processing the Results:
+    1. (a)
+      Ultimately, each token $t_{t}$ in Bloom will have a corresponding logits distribution matrix $logit_{t}$.
+    2. (b)
+      These logits can be directly used for subsequent training in the Bloom model.
+Appendix C Computation and Communication Complexity
 One of the key advantages of FedMKT is its computational efficiency.
@@
 This approach reduces communication overhead, allowing for more efficient data transmission and enhancing scalability in federated learning scenarios.
-Overall, the computational efficiency and reduced communication complexity of FedMKT make it a promising framework for federated mutual knowledge transfer for large and small Language models, particularly in scenarios where computational resources and communication bandwidth are limited.
-Setting Server Client-1 Client-2 Client-3 Client-4
-Heterogeneous LLaMa2- 7B GPT-2-xlarge(1.5B) OPT-1.3B Bloom-1.1B LLaMa2-1.3B
-Homogeneous LLaMa2- 7B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B
-Homogeneous LLaMa2- 7B OPT-1.3B OPT-1.3B OPT-1.3B OPT-1.3B
-One-to-One LLaMa2- 7B - - - LLaMa2-1.3B
-One-to-One LLaMa2- 7B - OPT-1.3B - -
-Table 1: FedMKT: LLM and SLMs Setting
-Task Method GPT-2-xlarge OPT-1.3B Bloom-1.3B LLaMa2-1.3B
-RTE SLM-ZS 52.4 52.7 52.7 49.8
-Standalone 65.7 62.5 58.1 55.6
-FedMKT-SLM 70.4 65.7 61.7 58.8
-WIC SLM-ZS 49.8 50.8 50 50
-Standalone 59.3 52.1 59.1 50.7
-FedMKT-SLM 63.2 62.2 61.1 51.9
-BoolQ SLM-ZS 61.3 58.4 59.0 61.0
-Standalone 71.1 74.1 69.9 69.7
-FedMKT-SLM 75.1 76.8 71.4 75.1
-CQA SLM-ZS 36.7 41.9 33.8 30.1
-Standalone 56.0 58.6 44.7 56.7
-FedMKT-SLM 58.3 60.5 50.8 57.0
-OpenBookQA SLM-ZS 22.4 23.4 19.6 23.2
-Standalone 27.4 26.6 21.0 25.8
-FedMKT-SLM 27.8 28.0 23.2 27.2
-ARC-E SLM-ZS 58.3 57.0 51.5 53.1
-Standalone 59.3 57.9 56.9 60.3
-FedMKT-SLM 59.8 59.6 57.5 60.8
-ARC-C SLM-ZS 25.0 23.4 23.6 26.7
-Standalone 28.2 28.4 24.9 28.5
-FedMKT-SLM 30.2 29.4 26.6 30.0
-S-NI SLM-ZS 5.0 5.2 5.1 5.8
-Standalone 27.9 26.1 10.6 33.4
-FedMKT-SLM 34.2 36.0 15.1 37.3
-DialogSum SLM-ZS 5.4 6.4 4.9 5.7
-Standalone 22.3 19.8 13.1 21.4
-FedMKT-SLM 23.2 20.9 14.9 21.6
-Table 2: Heterogeneous Setting SLMs Performance Comparison: FedMKT versus Baseline.
-4 Experiments
-4.1 Setup
-We set up a scenario involving three clients and one server to evaluate the FedMKT framework using various LLMs and SLMs.
-Models.
-We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al.
-(2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al.
-(2019), OPT-1.3B Zhang et al.
-(2022a), Bloom-1.1B Scao et al.
-(2022) and LLaMa2-1.3B Xia et al.
+Appendix D More on Experimental Details
+D.1 Hyperparameter Settings
+LoRA Parameters. We utilized the PEFTMangrulkar et al. (2022) library with the following configurations: r=8, lora_alpha=16, lora_dropout=0.05.
+Common Parameters for LLM and SLMs.
+We set batch_size=4, used the AdamW optimizer with adam_beta1=0.9 and adam_beta2=0.95.
+The warmup_ratio was set to 0.008, the weight_decay was 0.1, max_grad_norm was 1.0.
+The $\lambda$ was 0.9.
+The number of training rounds for all data is within 10 and the number of training rounds for different datasets may be different.
+LLM Parameters.
+During distillation, the local epoch R was set to 1.
+The learning rates $\eta_{\omega}$ were specified as 3e-5 for the datasets RTE/WIC/BoolQ/CQA/ARC-C/DialogSum/S-NI, and 2e-5 for ARC-E.
+SLM Parameters.
+During training for the four clients, the local epoch E was set to 1.
+The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3b", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1b1", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3b", the same learning rates as for the LLM were used.
+D.2 Data Splitting
+For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum, we randomly split the training data into five equal parts, with one part serving as the public dataset and the remaining four parts as private dataset for the four clients.
+All these datasets(including train, validate, test) were downloaded from HuggingFaceLhoest et al.
 (2021).
-In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One.
-Table 1 details the model configurations for both the LLM and SLMs in the distinct scenarios.
-Datasets.
-We comprehensively evaluate FedMKT on 7 QA datasets and 2 instruction-following datasets.
-Specifically, for QA tasks, we use RTE Wang et al.
-(2019), WTC Wang et al.
-(2019), BoolQ Clark et al.
-(2019), CommonsenseQA(CQA) Talmor et al.
-(2018), OpenBookQA Mihaylov et al.
-(2018), ARC-E Clark et al.
-(2018), ARC-CClark et al.
-(2018) to evaluate FedMKT.
-As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al.
-(2022), DialogSum Chen et al.
-(2021).
-Baselines. We conduct a comparative analysis of our FedMKT framework against several baselines to evaluate its performance. These baselines included:
-  • LLM-ZS, which represents the zero-shot (ZS) capabilities of the large language model (LLM) on the server;
-  • SLM-ZS, which represents the zero-shot performance of the small language model (SLM) on the clients;
-  • Standalone, where each client independently fine-tunes its local model using its own private dataset;
-  • FedAvg, which represents the standard federated averaging algorithm and is only used in the SLM comparisons of homogeneous settings.
-  • LLM2SLM, where the LLM is not updated in the FedMKT and is LLM only used to transfer knowledge to SLM in the FedMKT. It is only used in the SLM comparisons of the One-to-One setting.
-  • Centralized-LLM, where the server’s LLM is fine-tuned locally using the entirety of the private datasets combined with a public distillation dataset.
-Evaluation Metrics.
-We evaluate the performance of fine-tuned LLM and SLMs on both QA datasets and instruction-following datasets.
-For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
-This approach allows us to comprehensively assess the models’ capabilities across different tasks, providing a clear comparison of their performance.
-4.2 Heterogeneous Setting: Performance Evaluations
-Tables 2 and 3 demonstrate the performance comparisons of our approach against other baselines in the heterogeneous setting.
-Specifically, Table 2 focuses on the performance of SLMs, while Table 3 highlights the performance of LLM.
-As illustrated in Table 2, FedMKT-SLM demonstrates notably superior performance in comparison to both SLM-ZS and Standalone models.
-For instance, in the RTE QA dataset, FedMKT-SLM demonstrates significant performance gains compared to both SLM-ZS and Standalone models across various SLMs.
-Specifically, when using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 34% and Standalone by 7%.
-Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 25% and Standalone by 5%.
-In the Bloom-1.3B SLM, FedMKT-SLM achieves a 17% improvement over SLM-ZS and a 6% advantage over Standalone.
-Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 18% and Standalone by 6%.
-In the DialogSum instruction-following dataset, FedMKT-SLM again demonstrates impressive results.
-When using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 330% and Standalone by 4%.
-Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 227% and Standalone by 6%.
-In the Bloom-1.3B SLM, FedMKT-SLM achieves a 204% improvement over SLM-ZS and a 14% advantage over Standalone.
-Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 279% and Standalone by 0.9%.
-This highlights that within our framework, SLMs are able to effectively leverage the knowledge transferred from LLM, leading to enhanced model capabilities.
-As illustrated in Table 3, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM.
-For instance, in the RTE QA dataset, FedMKT-LLM outperforms LLM-ZS by 30% and achieves a performance level that is nearly on par with Centralized-LLM, reaching approximately 96% of its fine-tuning performance.
-Similarly, in the DialogSum instruction-following dataset, FedMKT-LLM outperforms LLM-ZS by a significant 185% and achieves approximately 90% of the fine-tuning performance of Centralized-LLM.
-This significant achievement signifies that FedMKT-LLM effectively facilitates the acquisition of knowledge from all clients by the server.
-This process enables the LLM to be fine-tuned utilizing client data without necessitating access to the entire model.
-Moreover, the clients’ raw private data is not required to be uploaded to the LLM server, thus preserving privacy while enabling the LLM to leverage the collective knowledge of all clients.
-Task Method LLaMa2-7B
-RTE LLM-ZS 63.1
-Centralized-LLM 85.9
-FedMKT-LLM 82.3
-WIC LLM-ZS 50.3
-Centralized-LLM 70.4
-FedMKT-LLM 61.3
-BoolQ LLM-ZS 70.1
-Centralized-LLM 87.6
-FedMKT-LLM 84.6
-CQA LLM-ZS 39.5
-Centralized-LLM 69.5
-FedMKT-LLM 71.8
-OpenBookQA LLM-FS 31.8
-Centralized-LLM 39.2
-FedMKT-LLM 37.8
-ARC-E LLM-ZS 69.3
-Centralized-LLM 76.9
-FedMKT-LLM 76.1
-ARC-C LLM-ZS 40.0
-Centralized-LLM 48.9
-FedMKT-LLM 44.7
-S-NI LLM-ZS 12.0
-Centralized-LLM 49.3
-FedMKT-LLM 41.4
-DialogSum LLM-ZS 8.5
-Centralized-LLM 27.7
-FedMKT-LLM 24.2
-Table 3: Heterogeneous Setting LLM Performance Comparison: FedMKT versus Baseline.
-4.3 Homogeneous Setting: Performance Evaluations
-In the homogeneous setting, we conduct experiments with two different model configurations, as shown in Table 1.
-These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B.
-It’s important to note that the number of clients is four, and within each client, all the SLMs is homogeneous.
-In this setting, to facilitate a comparison with FedAvg, we incorporated a FedAvg process involving all four clients within the FedMTK-SLM framework subsequent to mutual knowledge transfer.
-Furthermore, it’s worth noting that the standalone score here represents the average score obtained from the four clients working independently.
-Tables 4 demonstrate the performance comparisons of our approach against other baselines in the homogeneous setting.
-The performance comparisons encompass both LLM and SLMs.
-In this setting, FedMTK-SLM exhibits better performance than the Standalone and FedAvg approaches due to the assistance provided by LLM.
-Meanwhile, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM.
-Task Method LLaMa2 OPT
-CQA SLM-ZS 30.1 41.9
-Standalone 56.4 58.1
-FedAvg 56.4 58.6
-FedMKT-SLM-FedAvg 58.6 61.5
-LLM-ZS 39.5 39.5
-Centralized-LLM 69.5 69.5
-FedMKT-LLM-FedAvg 68.8 71.3
-ARC-C SLM-ZS 26.7 23.4
-Standalone 30.4 28.5
-FedAvg 29.7 28.6
-FedMKT-SLM-FedAvg 31.7 29.9
-LLM-ZS 40.0 40.0
-Centralized-LLM 49.4 49.4
-FedMKT-LLM-FedAvg 46.2 46.2
-ARC-E SLM-ZS 53.1 57.0
-Standalone 60.3 58.8
-FedAvg 60.6 59.0
-FedMKT-SLM-FedAvg 61.7 60.0
-LLM-ZS 69.3 69.3
-Centralized-LLM 75.5 76.9
-FedMKT-LLM-FedAvg 74.9 75.8
-Table 4: Homogeneous Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
-4.4 One-to-One Setting: Performance Evaluations
-In the One-to-One setting, we conduct experiments with two different model configurations, as shown in Table 1.
-These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B.
-Noth that both configurations involve only one client.
-Tables 5 demonstrate the performance comparisons of our approach against other baselines in the one-to-one setting.
-The performance comparisons encompass both LLM and SLMs.
-In this setting, FedMTK-SLM exhibits better performance compared to both the Standalone and LLM2SLM approaches.
-Compared to Standalone, FedMTK-SLM achieves superior results due to the assistance provided by LLM.
-When compared to LLM2SLM, FedMTK-SLM further improves performance by updating the LLM, which incorporates knowledge from client-side SLM.
-These findings underscore the benefits of federated mutual knowledge transfer between LLMs and SLMs.
-Meanwhile, compared to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses LLMs’ zero-shot performance and rivals Centralized-LLM’s fine-tuning performance.
-Task Method LLaMa2 OPT
-RTE SLM-ZS 49.8 52.7
-Standalone 55.6 62.5
-LLM2SLM 57.0 64.3
-FedMKT-SLM 61.0 63.2
-LLM-ZS 63.1 63.1
-Centralized-LLM 85.9 85.9
-FedMKT-LLM 82.3 79.1
-CQA SLM-ZS 30.1 41.9
-Standalone 56.7 58.6
-LLM2SLM 56.76 59.1
-FedMKT-SLM 56.84 60.7
-LLM-ZS 39.5 39.5
-Centralized-LLM 69.5 69.5
-FedMKT-LLM 69.0 71.0
-ARC-C SLM-ZS 26.7 23.4
-Standalone 30.3 28.8
-LLM2SLM 30.1 29.6
-FedMKT-SLM 30.8 30.4
-LLM-ZS 40.0 40.0
-Centralized-LLM 49.4 49.4
-FedMKT-LLM 45.9 45.8
-ARC-E SLM-ZS 53.1 57.0
-Standalone 57.0 57.9
-LLM2SLM 60.7 59.1
-FedMKT-SLM 60.8 59.7
-LLM-ZS 69.3 69.3
-Centralized-LLM 75.5 76.1
-FedMKT-LLM 74.8 75.8
-Table 5: One-to-One Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
-5 Conclusions and Future Work
-In this work, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models.
-FedMKT bridges the gap between the server-side large language model (LLM) and clients’ small language models (SLMs), enabling bidirectional knowledge transfer while preserving data privacy.
-Through extensive experiments across three distinct scenarios — Heterogeneous, Homogeneous, and One-to-One — we have demonstrated the effectiveness of FedMKT By utilizing a range of publicly available LLMs and SLMs on various NLP text generation tasks, we’ve shown that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
-The results highlight the framework’s adaptability and practicality, paving the way for integrating LLMs into real-world applications while protecting data privacy.
-As large language models continue to transform the landscape of artificial intelligence, FedMKT offers a promising approach for their responsible and effective utilization.
-Limitations and Future works. Our current study is constrained by computational and storage resources, preventing us from exploring larger model sizes. We plan to address this in future research.
+For the S-NI dataset, we first processed the data using minillmGu et al.
+(2023) to retain samples with an output length greater than or equal to 11.
+From this processed data, we randomly selected 300 samples as the evaluation dataset.
+The remaining data was then split into five equal parts, with one part serving as the public dataset and the other four parts as private data for the four clients.
+D.3 Dataset Licenses
+For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum were downloaded from HuggingFaceLhoest et al.
+(2021) and under Apache License, Version 2.0.
+For the S-NI dataset, it was from minillmGu et al.
+(2023) and under MIT License.
+D.4 Machine Configuration
+The experiments were conducted on machines equipped with either 4 Nvidia V100 32G or 8 Nvidia V100 32G GPUs.

### v4

@@
 To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models.
-This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights.
+This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enhancing the LLM with clients’ unique domain insights.
 We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance.
-Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks.
+Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT by utilizing diverse public LLMs and SLMs on a variety of NLP text generation tasks.
 Empirical results demonstrate that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
+Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/fedmkt
 FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models
-Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
+Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 The Hong Kong University of Science and Technology, Hong Kong, China 2 WeBank Co., Ltd, Shenzhen, China Correspondence: tfanac@cse.ust.hk
 1 Introduction
-Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities.
-The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), and LLaMA Touvron et al.
+The emergence of Large Language Models (LLMs) has marked a revolutionary shift in artificial intelligence, significantly transforming our understanding of natural language processing capabilities.
+The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), Gemma2 Team et al.
+(2024), and LLaMa2 Touvron et al.
 (2023) with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks.
 Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al.
-(2023) Fan et al.
-(2023).
+(2023); Fan et al.
+(2023, 2024b).
 The primary challenges include domain-specific knowledge Privacy, constrained computing resources, and mutual knowledge transfer between the LLM and SLMs.
@@
 Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
-To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models.
+To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to improve the performance of both large and small language models.
 By leveraging the complementary strengths of federated learning and knowledge distillation, FedMKT facilitates effective mutual knowledge transfer between clients’ SLMs and the LLM owned by the server.
@@
 During each round of federated learning, the clients transmit the output logits of their updated SLMs on the public dataset to the server.
-The server then selectively aggregates and distills the knowledge encoded within these SLMs output logits into the server-side LLM.
+Subsequently, the server selectively aggregates and extracts the knowledge encoded within these SLMs output logits into the server-side LLM.
 This process allows the server LLM to incorporate the domain-specific knowledge learned by the clients, thereby enhancing its comprehensive capabilities.
@@
 This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
+Our work is tailored for text generation tasks within the context of LLMs and supports both Heterogeneous and Homogeneous scenarios between client SLMs.
+To our best knowledge, our work is the first published research in this field.
   • Selective Knowledge Transfer and Token Alignment.
@@
   • Empirical Evaluation and Performance Enhancement.
-Extensive experiments conducted based on various publicly available LLMs and SLMs demonstrate the competitive performance of FedMKT across a wide range of NLP text-generation tasks.
+Extensive experiments conducted based on various publicly available LLMs and SLMs, have shown that FedMKT exhibits competitive performance across a broad spectrum of NLP text-generation tasks.
 We evaluate FedMKT with Heterogeneous, Homogeneous, and One-to-One settings.
@@
 2.1 Model Heterogeneous Federated Learning
-Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning.
+Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning (FL) Yang et al.
+(2019); McMahan et al.
+(2017); Liu et al.
+(2021); Cheng et al.
+(2021); Fan et al.
+(2024a).
 Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures.
-Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task.
+Several methods have been introduced to accommodate clients with different model architectures participating in a federated learning task.
 These methods typically involve techniques such as knowledge distillation Hinton et al.
@@
 (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
-While previous works have mainly focused on computer vision scenarios, the literature has limitedly explored MHFL in LLMs.
-This gap motivates this study, which aims to explore MHFL in the context of LLMs.
 2.2 Federated Learning for LLMs
@@
 (2022), and FATE-LLM  Fan et al.
-(2023).
-For example, the FedPETuning Zhang et al.
-(2022b) has demonstrated a significant reduction in communication overhead, reducing 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting.
-These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods.
-These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement.
+(2023, 2024c).
+These findings indicate that FL clients, especially those with constrained computing and storage resources such as certain devices, can significantly profit from from PEFT approaches.
+These techniques facilitate the sharing of LLMs across diverse tasks, while necessitating the storage and updating of only a small number of parameters for each task, thereby reducing the overall computational and storage requirements.
 By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
@@
 The FedMKT primarily comprises two key modules: Bidirectional Token Alignment and Selective Mutual Knowledge Transfer.
-We will elaborate on these two modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
+We will elaborate on these two key modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
 3.1 Problem Definition
 We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$.
-Each client owns a local private dataset $\mathcal{D}_{k}$ with $N$ training samples, and all clients and the server share a public dataset $\mathcal{D}_{p}$.
+Each client owns a local private dataset denoted as $\mathcal{D}_{k}$ containing $N$ training samples, and all clients and server have access to a shared public dataset $\mathcal{D}_{p}$.
 The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without disclosing any private data.
@@
 FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner.
-We illustrate the workflow of FedMKT in Figure 1 and elaborate on the associated training algorithm in Algorithm 1.
+We illustrate the workflow of FedMKT in Figure 1 and Appendix A, elaborate on the associated training algorithm in Algorithm 1.
 Algorithm 1 FedMKT
@@
 When multiple LLaMa2 tokens align to a single Bloom token (e.g., ’util’ and ’ize’ aligning to ’utilize’), we handle them according to the mapping table.
-Please refer to Appendix B for more details.
+Please see Appendix B for further details.
 In FedMKT, a bidirectional token alignment process occurs before knowledge transfer between LLMs and SLMs.
@@
 $\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}_{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
-where $\ell_{\text{CE}}$ is the cross-entropy loss; the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
+where $\ell_{\text{CE}}$ represents the cross-entropy loss, and the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
 Next, the server and clients conduct selective knowledge transfer to each other.
 The motivation for applying selective knowledge transfer is that some clients’ knowledge may adversely affect the performance of LLM on the server and vice versa in a heterogeneous environment.
-Therefore, it is critical to guarantee that the knowledge transferred between the server and clients is positive to the performance of LLM and SLMs.
-To this end, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
+Therefore, it is critical to guarantee that the knowledge exchanged between the server and clients is positive to the performance of LLM and SLMs.
+To achieve this goal, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
 DualMinCE aims to select knowledge that is positive to the performance of the server’s LLM from clients and vice versa.
@@
 $\mathcal{L}_{3}$ $=\frac{1}{K}\sum_{k=1}^{K}(\lambda\mathcal{L}_{\text{FT}}^{g}+(1-\lambda)\mathcal{L}_{\text{KD}}^{g})$ (10)
-where $\lambda$ is the hyperparameter that controls the weight of mutual knowledge transfer.
+where $\lambda$ is the hyperparameter that regulates the significance of mutual knowledge transfer.
 Setting Server Client-1 Client-2 Client-3 Client-4
@@
 Datasets.
-We evaluate FedMKT comprehensively on 6 QA datasets and 2 instruction-following datasets.
+We evaluate FedMKT on 6 QA datasets and 2 instruction-following datasets.
 Specifically, for QA tasks, we use RTE Wang et al.
@@
 (2019), BoolQ Clark et al.
-(2019), CommonsenseQA(CQA) Talmor et al.
-(2018), ARC-E Clark et al.
-(2018), ARC-C Clark et al.
+(2019), CQA  Talmor et al.
+(2018), ARC-E and ARC-C Clark et al.
 (2018) to evaluate FedMKT.
@@
   • Centralized, in which the server’s LLM is fine-tuned locally using the datasets combining private datasets of involved clients and the public dataset.
-In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are used to fine-tune the LLM;
+In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are utilized to fine-tune the LLM;
   • Zero-Shot, representing the zero-shot capabilities of LLM or SLMs (without fine-tuning);
-  • Standalone, in which each client independently fine-tunes its local SLM using its private dataset;
-  • FedAvg, representing the standard federated averaging algorithm. FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
+  • Standalone, where each client fine-tunes its own SLM independently using its private dataset;
+  • FedAvgMcMahan et al.
+(2017), representing the standard federated averaging algorithm.
+FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
   • LLM2SLM, representing FedMKT involving one server with an LLM and one client with an SLM.
@@
 LLM2SLM is only used in the One-to-One setting.
-Evaluation Metrics. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
+Evaluation Metrics.
+For the QA datasets, we primarily use Accuracy as the metric for evaluation, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
+It’s worth noting that in our experiments, all methods across the three scenarios undergo zero-shot evaluation, and we utilize the lm-evaluation-harness packageGao et al.
+(2023) for evaluation purposes.
 4.2 Evaluation on Heterogeneous Setting
@@
 Tables 2 show that FedMKT performs superior over Zero-Shot and Standalone on all clients’ SLMs.
-Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM.
+Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% in relative terms on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM.
 These empirical results demonstrate that, by leveraging FedMKT, SLMs are able to effectively leverage the knowledge transferred from the LLM, leading to enhanced model capabilities.
@@
 The bottom sub-table of Table 3 shows that FedMTK performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM.
-For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
+For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% in relative terms on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
 Task Method S1: Server LLaMa2-7B S2: Server LLaMa2-7B
@@
 5 Conclusions
-In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models.
-FedMKT bridges the gap between the server-side LLM and clients’ SLM, enabling selective mutual knowledge transfer while preserving data privacy.
+In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for LLMs and SLMs.
+FedMKT bridges the gap between the server-side LLMs and clients’ SLMs, enabling selective mutual knowledge transfer while preserving data privacy.
 Through extensive experiments across three distinct scenarios, we have demonstrated that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
@@
 This issue warrants further investigation.
+Additionally, the study does not address the potential presence of spammer Zhu et al.
+(2012) among clients, which could negatively impact the performance of LLM in the server and the SLMs of other clients.
+Detecting spammers in the FedMKT setting is identified as an important research direction.
 Furthermore, our study is limited by computational and storage constraints, which preclude the exploration of larger language models.
@@
     2. (b)
-      Use edit distance as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary.
+      Use minimum edit distance(MinED) as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary.
+The recursion function for MinED in dynamic programming is elaborated in Fu et al.
+(2023).
     3. (c)
@@
 In scenarios where multiple tokens align to one, like the 2-to-1 case of ’util’ and ’ize’ mapping to ’utilize’, we consider ’utilize’ as a match for ’util’ based on an optimal vocabulary mapping.
+Figure 2 illustrates an example of token alignment between LLaMa2 and Bloom.
   3. 3.
@@
       These logits can be directly used for subsequent training in the Bloom model.
+Figure 2: An example for token alignment via MinED.
 Appendix C Computation and Communication Complexity
@@
 During training for the four clients, the local epoch E was set to 1.
-The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3b", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1b1", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3b", the same learning rates as for the LLM were used.
+The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3B", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1.1B", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3B", the same learning rates as for the LLM were used.
 D.2 Data Splitting
@@
 The remaining data was then split into five equal parts, with one part serving as the public dataset and the other four parts as private data for the four clients.
-D.3 Dataset Licenses
-For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum were downloaded from HuggingFaceLhoest et al.
-(2021) and under Apache License, Version 2.0.
-For the S-NI dataset, it was from minillmGu et al.
-(2023) and under MIT License.
-D.4 Machine Configuration
+D.3 Machine Configuration
 The experiments were conducted on machines equipped with either 4 Nvidia V100 32G or 8 Nvidia V100 32G GPUs.